OpenStudy (anonymous):

Which line has an x-intercept of 4?

OpenStudy (anonymous):

a.4x + y = 8 b.-4x + y = 8 c.2x+ 2y = -8 d.-2x + 2y = -8

OpenStudy (zale101):

if the x-intercept is 4, then there's a point given which is (4,0)

OpenStudy (anonymous):

so the answer is a???

OpenStudy (anonymous):

is the answer a

OpenStudy (anonymous):

X=4 and Y=0 plug does values to each equation and see which one gives you both sides being equal.